A wheel of radius 20 cm is pushed to move it on a rough horizontal surface. It is found to move through
a distance of 60 cm on the road during the time it completes one revolution about the centre. Assume
that the linear and the angular accelerations are uniform. The frictional force acting on the wheel by the
surface is -

Grade:

1 Answers

Sourabh Kumar
6 years ago
VCM=60/T,  W(omega)=2pie/T. On eliminating T from both the eqn we get V=30W/pie. 
Now velocity at point of contact “Vcontact” =W×R =(2pie/T)×20 =40pie/T=125.67/T  ’m/s’,(leftward)      since radius is 20cm
Vcontact > VCM
Therefor skidding takes place leftward,hence friction act along direction of Vcm